OpenStudy (tvinsant):

PLEASE HELP. I WILL MEDAL AND FAN. Solve the quadratic equation by completing the square -3x^2+9x=1 a. 3/2±√6/6 b. -3±√69/3 c. 3±√-3/3 d. 3/2±√69/6

OpenStudy (anonymous):

1. Factor by Grouping. 2. Complete the Square. 4. 3. Use the Quadratic. 4 ... irst set of parentheses,place each term to the first power: 9x 1 .

OpenStudy (aum):

-3x^2 + 9x = 1 divide throughout by -3: x^2 - 3x = -1/3 Add (b/2)^2 to both sides. b = -3, (b/2)^2 = (-3/2)^2 = 9/4: x^2 - 3x + 9/4 = -1/3 + 9/4 (x - 3/2)^2 = -4/12 + 27/12 = (27-4)/12 = 23/12 finish the rest of the steps...

OpenStudy (tvinsant):

sorry, could you walk me through it?

OpenStudy (aum):

\[ -3x^2 + 9x = 1 \\ \text{Divide throughout by -3:}\\ x^2 - 3x = -\frac 13 \\ \text{To complete the square, add } \left(\frac b2\right )^2 \text{ to both sides:}\\ b = -3,~~ \left(\frac b2\right )^2 = \left(\frac {-3}{2}\right )^2 = \frac 94 \\ x^2 - 3x + \frac 94 = -\frac 13 + \frac 94 \\ (x - \frac 32)^2 = -\frac {4}{12} + \frac {27}{12} = \frac {27-4}{12} = \frac{23}{12}\\ \text{Take square root on both sides:} \\ x - \frac 32 = \pm \sqrt{\frac {23}{12}} \\ x = \frac 32 \pm \sqrt{\frac {23}{12}} \\ \]Can you simplify the right hand side?

OpenStudy (tvinsant):

so the answer would be a. x=3/2±√6/6 ?

OpenStudy (tvinsant):

lol i don't know how you could simplify the equation any further….

OpenStudy (tvinsant):

for sure it's either a. or d.

OpenStudy (aum):

The right hand side has a radical in the denominator. So you have to rationalize the denominator. \[ x = \frac 32 \pm \sqrt{\frac {23}{12}} \\ x = \frac 32 \pm \sqrt{\frac {23*3}{12*3}} = \frac 32 \pm \frac{\sqrt{69}}{6} \]

OpenStudy (tvinsant):

oh okay. thank you so much!